# SIV-Bench: A Video Benchmark for Social Interaction Understanding and Reasoning

### v1

@@ Abstract
The rich and multifaceted nature of human social interaction, encompassing multimodal cues, unobservable relations and mental states, and dynamical behavior, presents a formidable challenge for artificial intelligence. To advance research in this area, we introduce SIV-Bench, a novel video benchmark for rigorously evaluating the capabilities of Multimodal Large Language Models (MLLMs) across Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP). SIV-Bench features 2,792 video clips and 8,792 meticulously generated question-answer pairs derived from a human-LLM collaborative pipeline. It is originally collected from TikTok and YouTube, covering a wide range of video genres, presentation styles, and linguistic and cultural backgrounds. It also includes a dedicated setup for analyzing the impact of different textual cues—original on-screen text, added dialogue, or no text. Our comprehensive experiments on leading MLLMs reveal that while models adeptly handle SSU, they significantly struggle with SSR and SDP, where Relation Inference (RI) is an acute bottleneck, as further examined in our analysis. Our study also confirms the critical role of transcribed dialogue in aiding comprehension of complex social interactions. By systematically identifying current MLLMs’ strengths and limitations, SIV-Bench offers crucial insights to steer the development of more socially intelligent AI. The dataset and code are available at https://kfq20.github.io/sivbench/.
@@ 1 Introduction
The development of Multimodal Large Language Models (MLLMs), which can process text, images, and video, marks a significant step toward more human-like AI [46, 53, 64, 20, 55]. These models have shown strong performance across a range of tasks, driving progress in areas such as visual reasoning, video captioning, and multimodal dialogue. As these capabilities expand, there is a growing need for benchmarks that can evaluate model performance, uncover limitations, and guide future research [13, 11, 61, 37, 19]. One critical yet underexplored area is the understanding and reasoning about social interactions — a core aspect of social intelligence that encompasses not only observable behaviors but also implicit mental states and social relationships governing behaviors such as forming bonds, exchanging information, and coordinating actions [4, 42]. However, current video benchmarks—whether designed for specific tasks like object segmentation [17, 54], captioning [36], and activity recognition [60], or for broader video understanding across various domains [24, 25]—still lack a detailed focus on social interactions.
@@
We define the capacity of MLLMs to understand and reason social interactions in video through three core, interrelated dimensions: 1) Social Scene Understanding (SSU) is foundational, enabling the recognition of visible elements such as objects, environments, and socially salient human features like body movements, clothing, and physical appearance. Without reliable scene perception, models cannot anchor their interpretations in contextually relevant cues. 2) Social State Reasoning (SSR) is essential for interpreting the unobservable states of interaction, such as emotions, intents, attitudes, and interpersonal relationships, which guide and shape behavior [44, 52]. This capacity allows models to move beyond surface-level features and grasp the underlying states. 3) Social Dynamics Prediction (SDP) enables the model to reason about how interactions proceed over time or under alternative conditions, capabilities essential for a flexible and human-like understanding of social scenarios [39, 6]. It involves factual prediction and counterfactual prediction. The former anticipates upcoming actions or emotional shifts. The latter examines how changes in social relations or behaviors might alter the outcome of an interaction.
In this work, we introduce SIV-Bench (Social Interaction Video Benchmark), a novel benchmark designed to evaluate MLLMs across the above three core dimensions. SIV-Bench is fundamentally organized around social relationships, recognizing their critical role in shaping social interaction [47, 5, 16]. Specifically, SIV-Bench is built on Fiske’s Relational Models Theory [12], categorizing social interactions via four foundational models (Communal Sharing, Authority Ranking, Equality Matching, and Market Pricing), which are instantiated through 14 specific relation types (e.g., parent-child, friends, colleagues). This relational schema underpins fine-grained evaluation across three key dimensions. Understanding relational context is vital for SSU, as the salience and interpretation of actions, expressions, and environmental cues often depend on the relationship (e.g., a gaze between colleagues versus lovers). It is also central to SSR, where mental state inferences are modulated by relational roles (e.g., a critical comment from a mentor versus a stranger implies different states). For SDP, predictions of future behaviors and counterfactual reasoning are shaped by relational norms and history (e.g., conflict resolution strategies differ between siblings and business rivals).
SIV-Bench comprises 2,792 video clips sourced from TikTok and YouTube. Its video corpus ensures robust evaluation by encompassing diverse genres (such as daily life recordings, movie clips, sports footage, and animation), varied presentation styles (including first-person, third-person, phone calls, and solo multi-role sketches) and content reflecting multiple linguistic and cultural backgrounds (like English, Japanese, and Spanish). A dedicated question-answer (QA) pipeline, combining large language models with human verification, produces 8,728 high-quality multiple-choice questions (MCQs). These MCQs are designed to target specific facets within our three core assessment dimensions. Furthermore, to systematically assess the impact of linguistic cues, SIV-Bench incorporates audio tracks and provides videos under three distinct subtitle conditions: the original version (origin), a version with transcribed and translated dialogue added (+sub), and a version with all original on-screen text removed (-sub), as illustrated in Figure 4
In the experiments, we evaluate leading commercial(e.g., Gemini-2.5-Pro [9], o4-mini [33]), open-source (e.g., InternVL3 [63], Qwen2.5-VL [3]), and specialized video models (e.g., LLaVA-OneVision [23], LLaVA-Video [59]). The results show clear performance differences across the three dimensions. Models perform relatively well in SSU but struggle with SSR, where Relation Inference proves especially challenging. Common failure patterns include confusion between primary and secondary relationships, misleading contextual cues, weak commonsense reasoning, and missed perceptual details. SDP also poses difficulties, though top models handle counterfactuals relatively well. Our study also underscores the general importance of linguistic cues, as transcribed subtitles consistently aid overall comprehension while their removal typically hinders performance.
@@
Our key contributions are as follows: 1) We propose a novel analytical framework that structurally decomposes the complex task of multimodal social interaction understanding and reasoning into three core, interrelated dimensions, each further detailed into fine-grained sub-tasks. 2) We introduce SIV-Bench, a new video benchmark specifically curated for the analysis and comprehension of complex real-world social interactions. SIV-Bench comprises 2,792 real-world video clips representing 14 distinct social relationship types, and features 8,792 high-quality question-answer pairs generated through a human-LLM collaborative pipeline. 3) Our comprehensive experiments on diverse MLLMs reveal the limitations in their current capacity for deep human social understanding. These findings offer crucial insights to direct future research toward advancing artificial social intelligence.
@@ 2 SIV-Bench
This section details SIV-Bench, our novel video benchmark developed to evaluate MLLMs’ capabilities in understanding and reasoning within social interaction scenarios. Figure 1 offers some illustrative examples from SIV-Bench. The video corpus features interactions curated across 4 primary relational models, further detailed into 14 specific social relations, to ensure comprehensive coverage of real-world scenarios. The associated QA pairs are meticulously designed to probe model capabilities across 3 major task categories and 10 fine-grained aspects, targeting understanding, reasoning, and prediction in social contexts. Figure 2 shows the construction pipeline of SIV-Bench, including video collection (Section 2.1), QA generation and filtering (Section 2.2). Finally, to contextualize its contributions, SIV-Bench is also compared with other relevant benchmarks and datasets (Section 2.3).
Figure 2: The SIV-Bench construction pipeline, detailing the data collection process (left), and the QA generation & filtering process (right) with human-LLM collaboration. In the diagram,   blocks indicate content (like keywords, video and QA) generation steps,   blocks represent validation or modification stages, and   trapezoids signify a filtering and removal phase.
@@ 2.1 Video Collection
(a) Relation type distribution
(b) Video length distribution
(c) Language word-cloud
Figure 3: Video statistics for SIV-Bench: (a) Distribution of social relation types. (b) Distribution of video lengths, with an average of 32.49s. (c) Word cloud illustrating language diversity.
@@
Firstly, we utilize GPT-4o-mini [20] to generate comprehensive search keywords for each of the 14 relationship types (the word-clouds are shown in Figure 7), specifically including terms associated with varying degrees of intimacy, both positive (e.g., "love", "encourage") and negative (e.g., "conflict", "fight"). Leveraging these keywords, we conduct targeted searches and download initial video candidates from TikTok and YouTube platforms using Python libraries such as TikTokApi and yt-dlp, yielding approximately 5000 raw video clips. Each video is then manually reviewed by the authors to ensure it contains clearly observable and meaningful social interactions. Videos are excluded if they do not depict clear social interaction (e.g., a vlogger speaking directly to the camera without interacting with others), if the interaction context does not fit within a set of well-defined interpersonal scenarios (e.g., an interview setting with scripted dialogue), or if the dominant interaction is difficult to identify due to the presence of multiple overlapping social dynamics (e.g., a large multi-generational family posing for a group photo). These criteria are designed to ensure that each included video primarily features one interpretable and coherent type of interpersonal interaction, allowing for more consistent analysis of social behaviors across the dataset.
The final SIV-Bench comprises 2,792 curated video clips, with statistics shown in Figure 3. The collection showcases a rich diversity in social relationships (Figure 3(a)), where Communal Sharing interactions are the most represented category, with parent-child scenarios alone accounting for 18% of the videos. The other three relational models are also well represented, ensuring broad interpersonal coverage. In terms of duration (Figure 3(b)), clips average 32.49 seconds.
Figure 4: Illustration of the three subtitle conditions applied in SIV-Bench: ‘Origin’ (original video with existing on-screen text), ‘-Subtitle’ (original text removed), and ‘+Subtitle’ (transcribed and translated dialogue added).
While most fall in the 10–20 second range, the dataset spans a wide distribution, including many short clips (under 10 seconds) and a significant number over 60 seconds. Linguistically (Figure 3(c)), English serves as the predominant language, yet SIV-Bench incorporates a multicultural dimension by including content in various other languages such as Spanish, Filipino, Korean, and Japanese, thereby reflecting a greater diversity of real-world social scenarios [30, 48]. More details about video diversity (like genre and style) are shown in Appendix A.2.
@@
To evaluate how different forms of textual information affect MLLMs’ understanding of social interactions (Figure 4), we implement specific subtitle processing methods. Many original videos (‘Origin’) contain embedded on-screen text that often serves as scene descriptors or keywords (e.g., the ‘Buy and Sell’ text overlay). To focus on visual and auditory cues, we create a ‘-Subtitle’ version by removing such original textual overlays using video-subtitle-remover [56]. Conversely, to provide full access to spoken dialogue, we generate a ‘+Subtitle’ version. We employ Whisper-large-v3 [38] for audio transcription of the dialogue, and then use GPT to translate these transcriptions into English, ensuring consistent and high-quality subtitles (e.g., ‘He sold us his iPhone 12 with 256GB,’ as shown in ‘+Subtitle’).
@@ 2.2 QA Composing
Our Question-Answer (QA) composition process begins with generating an initial set of diverse QA pairs for each video, leveraging the capabilities of Gemini-2.0-Flash [15] which supports full video input. The prompts are carefully designed to elicit questions that span a broad range about social interactions, forming the basis for our SSU, SSR, and SDP evaluation dimensions. SSU focuses on descriptive observations of scenes, actions, and attributes; SSR targets inferences about intents, emotions, and relations; and SDP involves factual/counterfactual predictions. This stage typically yielded an average of 10 versatile QA pairs per video, designed to cover these core dimensions. Then we employ GPT-4o-mini to generate four distractors for each QA pair. These distractors are crafted to be contextually relevant to the question while remaining clearly distinguishable from the correct answer. We find that separating QA creation and distractor generation into two stages significantly improves quality and reduces ambiguity over generating all options simultaneously by a single model.
For quality assurance and to establish a gold standard, all generated QAs then undergo a rigorous consensus validation step [7, 8, 2]. We utilize three expert models (Gemini-2.0-Flash, Gemini-2.0-Pro, and GPT-4o-mini) to independently answer all questions. The QAs on which these models agree constitute our primary high-confidence set (Subset 1). The remaining QAs, where model agreement is not achieved, are flagged for further review or specialized handling (Subset 2). The following parts detail the subsequent processing strategies for each subset:
Subset 1: For Subset 1, which comprises QAs where expert models initially reached consensus (potentially indicating a need to increase challenge or verify video dependency), we implement a two-step filtering process. First, GPT-4o-mini is utilized to answer these questions without any video input. Questions that are answered correctly under this condition are removed. While this is a stringent criterion, as the model might occasionally guess correctly without sufficient information, it serves as an effective filter to ensure the questions necessitate video-based understanding, a worthwhile trade-off given our large initial pool of generated QAs. Second, the remaining questions are then scored for difficulty by GPT-4o-mini on a 1-to-5 scale. From this pool, we select the top 1-2 highest-scoring (i.e., most difficult) questions per video. Following these automated filtering stages, the authors conduct a final cross-validation review of these selected QAs to ensure their overall quality, relevance, and appropriate difficulty level. This multi-stage process ultimately yields 3,273 high-quality QAs.
Figure 5: Detailed statistics of Question-Answer pairs in SIV-Bench, showing the distribution across the 10 fine-grained sub-tasks.
Subset 2: We recruit a team of 20 human annotators to verify the non-consensus items. Each QA-pair is reviewed by at least two people. We only retain questions for which all reviewing annotators independently and unanimously select the same answer option, which is then established as the ground truth. This human validation process results in 3,096 QAs. Additionally, these annotators create novel and challenging questions along with corresponding answers. They follow specific instructions designed to target diverse abilities across the 10 fine-grained tasks. The questions are intentionally crafted to be difficult, requiring detailed understanding or complex reasoning. Appendix B.3 provides full details on our human annotation process, including demographic statistics on the 20 annotators, the guidelines for them, and the custom interface utilized. For these 2,359 newly human-generated QAs, plausible distractors are then efficiently created using GPT-4o-mini.
To ensure the quality and consistency of QA pairs, all curated QAs undergo a final automated refinement stage, where a GPT-4o-mini processes both correct answers and distractors. The goal is to standardize linguistic style and normalize option lengths. This minimizes superficial cues that models might exploit, improves test accuracy, and reduces variation from annotators or earlier generation steps. All the prompts and details in this section can be found in Appendix B.1 and B.2.
Finally, SIV-Bench contains a total of 8,728 QA pairs. Their distribution across core assessment dimensions and fine-grained tasks is illustrated in Figure 5. SRR accounts for the largest share, with 4,951 QAs, highlighting the benchmark’s focus on reasoning in social interactions. This is followed by SDP with 2,664 QAs, and SSU with 1,113 QAs. Within these categories, sub-tasks like Intent Inference (2,132 QAs) and Factual Prediction (1,449 QAs) show especially high concentrations, supporting a rich evaluation of MLLM capabilities. A detailed statistical analysis of the SIV-Bench QA pairs, confirming their structural balance and diversity (e.g., option length, answer distribution, question complexity, interrogative types), is provided in Appendix B.4 with Figure 17.
2.3 Comparison with Previous Datasets & Benchmarks
Table 1 highlights the key differences between SIV-Bench and previous datasets and benchmarks. The earlier datasets that approach social interaction primarily through the lens of social relations, such as DialogRE [58] (text-only), PIPA [45] (image-only), and ViSR [26], are often restricted to the singular task of relation recognition or limited by single modalities. They also tend to lack fine-grained task diversity and face scalability challenges due to manual annotation.

### v2

@@ Abstract
Understanding social interaction, which encompasses perceiving numerous and subtle multimodal cues, inferring unobservable mental states and relations, and dynamically predicting others’ behavior, is the foundation for achieving human-machine interaction. Despite rapid advances in Multimodal Large Language Models (MLLMs), the rich and multifaceted nature of social interaction has hindered the development of benchmarks that holistically evaluate and guide their social interaction abilities. Based on social relation theory, which has been widely regarded as a foundational framework for understanding social behavior, we provide SIV-Bench, a novel video benchmark for systematically evaluating MLLMs’ capabilities across Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP). SIV-Bench features 2,792 originally collected video clips and 5,455 meticulously generated question-answer pairs derived from a human-LLM collaborative pipeline. It covers 14 typical relationships, diverse video lengths, genres, presentation styles, and linguistic and cultural backgrounds. Our comprehensive experiments show that leading MLLMs perform relatively well on SSU but remain weak on SSR and SDP, with the systematic confusion in relation inference as a key bottleneck. An in-depth analysis of the reasoning process attributes MLLMs’ suboptimal performance to misalignment with human thoughts and insufficient reasoning depth. Moreover, we find audio and subtitles aid in reasoning-intensive SSR and SDP. Together, SIV-Bench offers a unified testbed to measure progress, expose limitations, and guide future research toward more socially intelligent MLLMs. We release the dataset and code at our project website: https://kfq20.github.io/sivbench/.
SIV-Bench: A Video Benchmark for Social Interaction Understanding and Reasoning
Fanqi Kong1,2, Weiqin Zu2,4, Xinyu Chen1, Yaodong Yang1, Song-Chun Zhu1,2,3, Xue Feng2thanks: Corresponding Author. 1Peking University   2State Key Laboratory of General Artificial Intelligence, BIGAI 3Tsinghua University   4ShanghaiTech University kfq20@stu.pku.edu.cn, fengxue@bigai.ai
@@ 1 Introduction
The rapid development of Multimodal Large Language Models (MLLMs) capable of processing text, images, and video has driven strong performance across tasks such as visual reasoning, video captioning, and multimodal dialogue (Team et al., 2023; Wu et al., 2024a; Zhu et al., 2025b; Hurst et al., 2024; Yang et al., 2024; Wu et al., 2026). As these capabilities expand, there is a growing need for benchmarks that can evaluate model performance, uncover limitations, and guide future research (Fu et al., 2024; Fang et al., 2024; Zhou et al., 2024; Qiang et al., 2025; Huang et al., 2024; Li et al., 2023; Xie et al., 2026). One critical yet underexplored area is the social interaction understanding and reasoning, which is a core aspect of social intelligence that encompasses not only observable behaviors but also implicit mental states and social relationships governing behaviors such as forming bonds, exchanging information, and coordinating actions (Berger et al., 1972; Smith-Lovin and Heise, 1988). However, existing video benchmarks, whether designed for specific tasks like video object segmentation (Ding et al., 2025; Athar et al., 2025), captioning (Chen et al., 2025; Wu et al., 2025), and fine-grained action understanding (Perrett et al., 2025), or for broader video understanding (Wang et al., 2025; Li et al., 2024b, c), still struggle to systematically probe MLLMs’ understanding of the multifaceted nature of social interaction.
@@
To address this gap, we first decompose the capacity to understand and reason about social interaction into three core, interrelated dimensions. 1) Social Scene Understanding (SSU) is foundational, enabling the recognition of visible elements such as objects, environments, and socially salient human features like body movements, clothing, and physical appearance. Reliable scene perception is required to ground interpretations in relevant cues. 2) Social State Reasoning (SSR) is essential for interpreting the unobservable states of interaction, such as emotions, intents, attitudes, and interpersonal relationships, which guide and shape behavior (Strachan et al., 2024; Wu et al., 2020; Kong et al., 2025). This capacity allows models to move beyond surface-level features and grasp the underlying states. 3) Social Dynamics Prediction (SDP) enables the model to reason about how interactions proceed over time or under alternative conditions, capabilities essential for a flexible and human-like understanding of social scenarios (Ramnani and Miall, 2004; Byrne, 2016). It involves factual prediction and counterfactual prediction. The former anticipates upcoming actions or changes in social states, while the latter examines how alterations in social scenes or states can affect interaction outcome.
Second, to systematically and comprehensively evaluate the three dimensions, we introduce SIV-Bench (Social Interaction Video Benchmark) grounded in social relation theory, which has been widely regarded as a foundational framework for understanding social behavior (Thibaut, 2017; Burkitt, 1997; Hartup, 1989). Specifically, SIV-Bench is built on Fiske’s Relational Models Theory (Fiske, 1992), categorizing social interactions via four foundational models (Communal Sharing, Authority Ranking, Equality Matching, and Market Pricing), instantiated through 14 relation types (e.g., parent-child, friends, colleagues). This relational context underpins all three dimensions. It conditions how cues are interpreted in SSU (e.g., a gaze between colleagues vs. lovers), modulates mental-state inference in SSR (e.g., criticism from a mentor vs. a stranger), and shapes SDP by constraining future behaviors and counterfactuals through relational norms and history (e.g., siblings vs. business rivals). SIV-Bench comprises 2,792 video clips sourced from TikTok and YouTube, and a QA pipeline that combines adversarial filtering with human verification produces 5,455 high-quality questions. To systematically assess the contribution of linguistic cues, SIV-Bench further provides audio tracks and three subtitle conditions: the original version (origin), a version with transcribed dialog added (+sub), and a version with all on-screen text removed (-sub).
Figure 2: The SIV-Bench construction pipeline, detailing the data collection process (left), and the QA generation & filtering process (right) with human-LLM collaboration. In the diagram,   blocks indicate content (like keywords, video and QA) generation steps,   blocks represent validation or modification stages.
In the experiments, we evaluate a broad set of models, including leading commercial MLLMs, strong open-source MLLMs, and video-specialized models. Overall, models are comparatively strong on SSU but weak on high-level reasoning, particularly within SSR, where relation inference exhibits systematic confusions. Our fine-grained analyses reveal recurring error sources, including selecting plausible but secondary relations, over-reliance on contextual cues, insufficient commonsense social reasoning, and missed perceptual details. SDP remains challenging as well, with counterfactual cases being relatively better handled by top models. We further study textual cues via controlled subtitle ablations and observe a task-dependent effect. Removing language typically has a limited impact on basic perception, but can degrade performance on inference-intensive SSR/SDP cases. Furthermore, on a hard diagnostic subset with brief explanations and a human baseline, we observe a large remaining gap to humans, underscoring the difficulty of robust, human-aligned social reasoning.
@@
Our key contributions are as follows: 1) We propose a novel analytical framework that structurally decomposes the complex task of multimodal social interaction understanding and reasoning into three core, interrelated dimensions, each further detailed into fine-grained sub-tasks. 2) We introduce SIV-Bench, a new video benchmark specifically curated for the analysis and comprehension of complex real-world social interactions. SIV-Bench comprises 2,792 real-world video clips representing 14 distinct social relationship types, and features 5,455 high-quality question-answer pairs generated through a human-LLM collaborative pipeline. 3) Our comprehensive experiments on diverse MLLMs reveal the limitations in their current capacity for deep human social understanding.
@@ 2 SIV-Bench
This section presents the construction of SIV-Bench, including video collection (Section 2.1), QA composing (Section 2.2), and a comparison with existing benchmarks (Section 2.3). Figure 1 offers some illustrative examples from SIV-Bench. Figure 2 shows the construction pipeline.
@@ 2.1 Video Collection
Firstly, we utilize GPT-4o-mini (Hurst et al., 2024) to generate comprehensive search keywords for each of the 14 relationship types (the word-clouds are shown in Figure 8), specifically including terms associated with varying degrees of intimacy, both positive (e.g., "love", "encourage") and negative (e.g., "conflict", "fight"). Leveraging these keywords, we conduct targeted searches and download initial video candidates from TikTok and YouTube platforms using Python libraries such as TikTokApi and yt-dlp, yielding approximately 5000 raw video clips. Each video is then manually reviewed by the authors to ensure it contains clearly observable and meaningful social interactions. Videos are excluded if they do not depict clear social interaction (e.g., a vlogger speaking directly to the camera without interacting with others), if the interaction context does not fit within a set of well-defined interpersonal scenarios (e.g., an interview setting with scripted dialogue), or if the dominant interaction is difficult to identify due to the presence of multiple overlapping social dynamics (e.g., a large multi-generational family posing for a group photo). These criteria are designed to ensure that each included video primarily features one interpretable and coherent type of interpersonal interaction, allowing for more consistent analysis.
The SIV-Bench comprises 2,792 curated video clips, with statistics shown in Figure 3. The collection showcases a rich diversity in social relationships (Figure 3(a)). Communal Sharing interactions are the most represented category, reflecting the naturalistic prevalence and psychological centrality of these relationships in daily life (Simão and Seibt, 2014; Kameda et al., 2005). The other three relational models are also well represented, ensuring broad interpersonal coverage. In addition to relationship diversity, the benchmark covers heterogeneous genres and filming/editing conventions, which helps test models under different visual narratives and interaction realizations. In terms of duration (Figure 3(b)), clips average 32.49 seconds. While most clips are 10–20 seconds, the dataset spans a wide distribution, including many short clips (under 10 seconds) and a significant number over 60 seconds. English predominates, but SIV-Bench also includes other languages, adding to multicultural diversity (detailed in Figure 9). Details on video diversity are in Appendix A.2.
(a) Relation type distribution.
(b) Video length distribution.
Figure 3: Video statistics for SIV-Bench.
@@
To evaluate how different forms of textual information affect MLLMs’ understanding of social interactions, we implement specific subtitle processing methods (Figure 4). Many original videos (‘Origin’) contain embedded on-screen text that often serves as scene descriptors or keywords (e.g., the ‘Buy and Sell’ text overlay). To focus on visual and auditory cues, we create a ‘-Subtitle’ version by removing such original textual overlays using video-subtitle-remover (YaoFANGUK, 2025). Conversely, to provide full access to spoken dialogue, we generate a ‘+Subtitle’ version. We employ Whisper-large-v3 (Radford et al., 2022) for audio transcription of the dialogue, and then use GPT to translate these transcriptions into English, ensuring consistent and high-quality subtitles (e.g., ‘He sold us his iPhone 12 with 256GB’).
Figure 4: Illustration of the three subtitle conditions.
Table 1: Comparison of various benchmarks, including total number of items (#Items, representing the number of videos, dialogues, images, etc.), number of QA pairs (#QAs), annotation method (Anno., M/A means manually/automatic manner), and tasks (Task Types: SU for Scene Understanding, SR for State Reasoning, DP for Dynamics Prediction). Note that Task Types here refer to general task categories, which include both social and physical scenarios. The table also shows whether each benchmark includes Multi-Person Interaction, covers Various Relations, is based on newly collected data (Original Collection), and provides subtitle/audio (S.A.).
Benchmark #Items #QAs Anno. Task Types Multi-Person Various Original S.A. SU SR DP Interaction Relations Collection Social Relation Inference DialogRE (Yu et al., 2020) 1,788 - M - PIPA (Sun et al., 2017) 37,107 - M - ViSR (Liu et al., 2019) 8,000 - M General Video Understanding and Reasoning VideoMME (Fu et al., 2024) 900 2,700 M MLVU (Zhou et al., 2024) 1,730 3,102 M VideoVista (Li et al., 2024c) 894 24,906 A Social-IQ 2.0 (Wilf et al., 2023) 1,000 6,000 M Social Genome (Mathur et al., 2025) 272 1,486 M Perception Test (Patraucean et al., 2023) 11,620 38,000 M MVBench (Li et al., 2024b) 3,641 4,000 A Video-Bench (Ning et al., 2023) 5,917 17,036 A&M EgoSchema (Mangalam et al., 2023) 5,063 5,063 A&M SIV-Bench 2,792 5,455 A&M
@@ 2.2 QA Composing
Our Question-Answer (QA) composition pipeline is designed to move beyond simple recognition tasks and ensure a high density of challenging reasoning problems. The process consists of three main stages: scalable generation, adversarial filtering, and human-in-the-loop curation.
@@
Scalable Generation. We begin by generating a large initial pool of diverse QA pairs for each video, leveraging the capabilities of Gemini-2.0-Flash with full video input. The prompts are carefully designed to elicit questions that span our SSU, SSR, and SDP evaluation dimensions. This stage typically yields an average of 10 QA pairs per video. Subsequently, we employ GPT-4o-mini to generate four distractors for each QA pair, ensuring they are contextually relevant yet clearly distinguishable from the correct answer. Separating distractor generation significantly reduces ambiguity compared to simultaneous generation.
Adversarial Filtering. To ensure SIV-Bench serves as a rigorous diagnostic tool rather than a saturated test set, we implement a strict Model-Consensus Filtering mechanism. We utilize three commercial models (Gemini-2.0-Flash, Gemini-2.0-Pro, and GPT-4o-mini) to independently answer all candidate questions. We adopt an adversarial approach: questions correctly answered by all models (indicating trivial reasoning or visual obviousness) are discarded. This can also prevent the model from choosing the correct answer without looking at the question itself or the video prompt (Li et al., 2025a). We focus exclusively on the remaining "hard" samples where model consensus is not achieved, as these disagreements signal the need for deeper social reasoning.
Figure 5: QA Distribution across 10 sub-tasks.
Human Verification and Authoring. The questions surviving the adversarial filter undergo a rigorous human-in-the-loop process to distinguish legitimate challenging samples from noise: (1) We recruit 20 human annotators to verify the non-consensus items. Each QA is reviewed by at least two annotators. Only questions where humans independently agree on the correct answer are retained. This yields 3,096 high-quality, model-challenging QAs. (2) To further expand the benchmark’s ceiling, annotators are instructed to generate novel, complex questions and their answers. These human-authored 2,359 QAs are designed to probe nuances often missed by model generators. Appendix B.2 details the annotation guidelines and interface. For all new human-generated QAs, distractors are created using GPT-4o-mini.
Finally, all curated QA pairs undergo an automated refinement stage to standardize linguistic style and option length, minimizing superficial cues that models might exploit. Full prompts are provided in Appendix B.1. Through this "Filter-then-Verify" pipeline, SIV-Bench ultimately comprises 5,455 high-quality QA pairs. The distribution across core dimensions is illustrated in Figure 5, featuring a strong emphasis on SSR and SDP, enabling rich evaluation of higher-order social intelligence. Appendix B.3 provides detailed statistics confirming structural balance and diversity.
2.3 Comparison with Existing Benchmarks